OpenStudy (anonymous):

A brush cost $2 more than a comb. The brush and comb together cost $3.78. Find the cost of each. I'm trying to use the equation 2x+x=3.78 but I don't think this is right. Help?

zepdrix (zepdrix):

so a brush cost 2 more than a comb ---> comb + 2 = brush brush and comb cost 3.78 -----------> brush + comb = 3.78 you have a system of equations, add them.. subtract them.. do something to where one of your variables cancel out and solve for the other =) then with the value you get, plug the value into one of your original equations to find the other! yayyyy math!

OpenStudy (anonymous):

@Carly33 what class is this? Have you done systems of 2 equations with 2 unknowns yet?

OpenStudy (anonymous):

Pre-Algebra.

OpenStudy (anonymous):

Ok... you often get these in pre-algebra classes where you haven't had the stuff zepdrix showed (it was right, just stuff you haven't seen yet)

OpenStudy (anonymous):

A good equation to use is "comb" + "2 dollars more than comb" = 3.78. Do you see why?

OpenStudy (anonymous):

so if comb is "x", you would have x + (2 + x) = 3.78 2x + 2 = 3.78 2x = 1.78 x = (1.78 / 2) = 0.89

OpenStudy (anonymous):

Ok, I think I just got it. x= brush & y=comb so 2 + 2y = 3.78 so y= .89??? and then I plug that back in, right?

OpenStudy (anonymous):

That's the comb price, 0.89. And the brush price is 2 dollars more, or 2.89. Check your work after, also... a very good habit: 0.89 + 2.89 = 3.78 And now you can say "yay math" :)

OpenStudy (anonymous):

Yay math, lol. thank you!!

OpenStudy (anonymous):

My point was that in pre-algebra, they don't teach you to use two different variables like x and y... they just take the short cut and want you to say "one variable" and "some number plus that variable". But the BEST way is to use the x and y... it's just assumed that doing that in this class at this point might be too much for you. :) You will get there soon enough though! :)

OpenStudy (anonymous):

Ok, thank you very much. Word problems just intimidate me.

OpenStudy (anonymous):

Be BRAVE :)

OpenStudy (anonymous):

And just take it slow... talk it out as you write the math expression to go with the word problem... it gets easier with practice.

OpenStudy (anonymous):

:) thanks again!

OpenStudy (anonymous):

you are most welcome :) Good luck!